Previous

Calendar

May

2007

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

29

JD119

30

JD120

1

JD121
Site Visit (SpB, JB): Details , Photos (16)

at 10:40 MST, removed 2nd PAR sensor + photodiodes from 26m

from 11:15-11:50 MST, adjusted the LI-6262 and LI-6251 span/N2 flow rates

Prius Rental

2

JD122
(Full "Flower" Moon)

3

JD123
Site Visit (SpB, JB, JH): Details , Photos (24)

Transported 5 new cylinders to the trailer

at 11:25 MST, changed the 0.5m inlet filter (for li-6251)

at 11:40 MST, swiped the kyrpton hyg

pre-cal on 2 co2-in-air cylinders, post-cal on 1 co2-in-air cylinder (in trailer)

Swapped 4 flasks at the west robot

Drove to C1!

Jeep Rental

4

JD124

5

JD125 (snow)
quacker system (hi-rate data) down from ~11:00-12:37 MST.

6

JD126

7

JD127

8

JD128 (sunny/clear)
Site Visit (SpB, JB, NT): Details , Photos (27)

from 12:20-12:45 MST, replaced domes on REBS Rnet sensor

re-aligned RM Young prop-vane at 25m, replaced prop-vane at 16m

very wet snow!

quacker system (hi-rate data) down from 16:32-20:26 MST.

Summer Jeep Rental (starts today)

9

JD129

10

JD130
Site Visit (SpB, JB, SO, KK): Details , Photos (21)

at 12:45 MST, added 2nd REBS Rnet sensor at 25m

Swapped 4 flasks at west robot

11

JD131
Site Visit (JH, SL, DK): Details

12

JD132

13

JD133
Mothers Day
JB to Durango

14

JD134

15

JD135 (rain/snow)
from 9-9:15 MST, the flow rate (ie, pump) for the LI-6262 dropped (temporarily).

16

JD136
at around 10:00 MST, the in-canopy csats dried out.

JB returns

17

JD137

18

JD138
Site Visit (SpB, JB, JH, KK): Details , Photos (18)

at 15:02 MST, russter2 shuts down (power outage, UPS has a problem)

at 16:38 MST, russter2 rebooted (UPS powered back up)

at 20:50 MST, russter2 shuts down again

at 21:02 MST, russter2 reboots itself

19

JD139
Site Visit (SpB, JB): Details , Photos (4)

at 12:49 MST, quacker shuts down (probably a lightning strike)

at ~19:30 MST, start running ops38

replaced UPS in the trailer with a borrowed UPS

20

JD140
at 7:42 MST, start running ops39

data system running with limited channels (and only 1-hz).

21

JD141
Site Visit (SpB, JB): Details , Photos (15)

at 10:15 MST, edgeport and network switch at tower replaced, quacker back up (all channels appear ok), now running ops40

from 10:50-11:28 MST, Changed profiler (li-6251) dessicant; (blow test from 11:21-11:28)



22

JD142

23

JD143

24

JD144 (cool but sunny)
Site Visit (JB, JH, DK, SLS): Details , Photos (9)

Swapped 4 flasks

3 inches of snow overnight

25

JD145

26

JD146

27

JD147

28

JD148
CU Holiday (Memorial Day)

29

JD149
between 11-12:30 MDT, an intense storm passed by the tower

the flow rate (ie, pump) for the LI-6262 was acting strange (on and off thoughout the day).

30

JD150
Site Visit (SpB, JB): Details , Photos (27)

at 12:30 MST, replaced regulator for span gas at tower

from ~11:30-12:30 MST, replaced ec gast pump with refurbished pump; made flow adjustments at LI-6262

at 11:50 MST, started running "met_02.csi" on MET logger (output RM Young in mV to quacker)

31

JD151 (blustery)
Site Visit (SpB, JB): Details , Photos (15)

from ~11:00-15:30 MST, replaced gelman inlet filters on profile system (1,2,6,10, and 21.5m)

from 10:52-12:30 MST, running ops41 (removed handar.9m & modified 2-d ati code)

at 12:37 MST, start running ops42 (removed 2-d ati, added back NCTC's)

Swapped 4 flasks at west robot (appears to be pump problems?)

at 10:45 MST, started running "met_03.csi" on MET logger (removed 9m handar output)

1

JD152

2

JD153